OpenStudy (sasogeek):

Integrate sin(2x + pi/6)dx from 0 to pi/6

OpenStudy (sasogeek):

\[\int\limits_{0}^{\pi/6}(2x + \pi/6)dx\]

OpenStudy (anonymous):

\[-\sqrt{3}/4\]

OpenStudy (sasogeek):

explain.... :/

OpenStudy (anonymous):

again check it...it can't be negative :/

OpenStudy (sasogeek):

...? I'm lost...how did u arrive at that answer?

OpenStudy (anonymous):

\[\sin(2x+\pi) = - \sin(2x)\] so, we can write it as -sin2x /6 then integrate

OpenStudy (anonymous):

i thinks thats wrong

OpenStudy (anonymous):

kishan:i think it is sin(2x+(pi/6))dx so let 2x+pi/6=t; differnentiaite 2dx=dt dx=dt/2 so it becomes integral of sin t dt/2 =-(cos t)/2 putting limits we get (cos 0 -cos pi/6)/2 =(1-1/2)/2 =1/4

OpenStudy (anonymous):

but cos(\[\pi\]/6)==\[\sqrt{3}\]/2

OpenStudy (anonymous):

oh sorry it will be (1-sqrt(3)/2)/2.